OpenStudy (anonymous):

f(x)=4^cosx + e^x , find f'(x)

OpenStudy (anonymous):

\[f'(x)= \ln(4) 4^{cosx} (-sinx) + e^x \]\[=-4^{cosx} \ln(4) sinx +e^x\]

OpenStudy (anonymous):

Can you show steps?

OpenStudy (anonymous):

\[\frac{d}{dx} A^{f(x)} = lnA \times A^{f(x)} \times f'(x)\]

OpenStudy (anonymous):

and e^x is the same

OpenStudy (anonymous):

Thanks!